OpenStudy (anonymous):

3(2x - 4) = 4(x + 1) solve for x

OpenStudy (anonymous):

\[6x-12=4x+4\] \[6x-4x=4+12\] \[2x=16\] \[x=8\]

OpenStudy (anonymous):

ohhh see i thought i was supposed to subtract 4 from 12... :/

OpenStudy (anonymous):

you can subtract 4 from -12 like this \[6x-12-4=4x\] \[-16=4x-6x\] \[-16=-2x\] \[x=8\]

OpenStudy (anonymous):

it's same just bit weird since you get minuses

OpenStudy (anonymous):

yeah... yeah i get it thank you very much :D

OpenStudy (anonymous):

6x-12=4x+4 2x=16 x=8